OpenStudy (anonymous):

A Chevy is traveling 10 mph faster than a Toyota. The Chevy is traveling at a rate of 60 mph. Let T = the the speed of the Toyota. Which formula below would correctly calculate the speed of the Toyota?

OpenStudy (jack1):

C = chevy speed = T + 10 = 60 T = toyota speed so if T + 10 = 60... T = ??